One year a couple invested a total of $10,000 in two projects. The first investment , a mini mall , made a 6% profit. The other profit investment, a skateboard park, made 12% profit. If their investments made $960 , how much invested at each rate?
Answer by ReadingBoosters(3246)

x+y = 10000
.06(x) + .12(y) = 960
Substitute into the second equation
.06(10000-y) + .12y = 960
600 - .06y + .12y = 960
600 + .06y = 960
.06y = 960 - 600
.06y = 360
y = $6,000 at 12%
x = 10000 - 6000 = $4,000 at 6%
Proof
.06(4000) + .12(6000)
240 + 720 = 960
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.